# thanks: $0$-0003-0728-2533thanks: $0$-0001-5501-6008

### v1

@@ Abstract
The internal structure and abundance of dark matter halos and subhalos are powerful probes of the nature of dark matter. In order to compare observations with dark matter models, accurate theoretical predictions of these quantities are needed. We present a fast and accurate method to describe the tidal evolution of subhalos within their parent halo, based on a semi-analytic approach. We first consider idealized N-body simulations of subhalos within their host halo, using a generalized mass density profile that describes their properties in a variety of dark matter models at infall, including popular warm, cold, and self-interacting ones. Using these simulations we construct tidal “tracks” for the evolution of subhalos based on their conditions at infall. Second, we use the results of these simulations to build semi-analytic models (SAMs) for tidal effects, including stripping and heating and implement them within the code galacticus. Our SAMs can accurately predict the tidal evolution of both cored and cuspy subhalos, including the bound mass and density profiles, providing a powerful and efficient tool for studying the post-infall properties of subhalos in different dark matter models.
@@ I Introduction
However, while the CDM model has shown excellent agreement with observations on large scales, such as the Cosmic Microwave Background and large-scale structure of galaxy distributions, reproducing some observations on subgalactic scales is challenging within the model [1, 2]. Examples of these challenges include the cusp-core problem [3, 4, 5], the too-big-to fail problem  [6], and diversity problem [7, 8]. To solve these issues, a number of alternative DM models have been proposed, including, e.g, warm dark matter [9, 10, 11, 12], fuzzy dark matter (FDM) [13, 14, 15, 16, 17, 18], self-interacting dark matter (SIDM) [19, 20], and primordial black holes [21, 22, 23].
A particularly powerful probe of the nature of DM is the abundance and internal structure of halos and subhalos. These are the bound hierarchical structures that form as a result of gravity, as the universe evolves. Alternate dark matter models predict different density profiles and abundance for small DM halos and subhalos. The demographics and internal structure of halos and subhalos already provide some of the most stringent limits on alternative dark matter models, including from observations of Milky Way satellites [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and strong gravitational lensing [34, 35, 29, 36, 37, 38, 39].
@@
One way to control numerical artifacts is to run idealized simulations, in which case only one subhalo is evolved in the host potential so that one can run high-resolution simulations with manageable computing resource. Using such methods, van den Bosch & Ogiya [48] carried out a detailed study of numerical artifacts in simulations resulting from gravitational softening, discreteness noise, and two-body relaxation. They also derived the requirements for obtaining properly converged results. However, in order to compute statistical properties of subhalos, one needs a model to describe the tidal evolution of subhalos so that one could look at the evolution of a population of subhalos within the context of a cosmological model. In order to achieve this, one can either build analytic models using appropriate approximations, or train non-parametric models such as the one presented in Ref. [49].
@@
In this work, we first run idealized N-body simulations to study the evolution of DM halos with different initial density profiles in a host gravitational potential, aiming to include a broad range of dark matter profiles that may be produced by different DM models. Previous studies [50, 51, 52, 53] have shown that, as subhalos evolve in the host, their maximum circular velocity, $V_{\mathrm{max}}$  [^1] The circular velocity is defined as $V=\sqrt{GM(r)/r}$., and the radius at which this maximum is reached, $R_{\mathrm{max}}$, follow a universal “tidal track” for a specific initial density profile. Both $V_{\rm max}/V_{\rm max,0}$ and $R_{\rm max}/R_{\rm max,0}$ are functions of fractional mass remaining in the subhalo and are not sensitive to how mass is stripped from the subhalos. We calculate the tidal tracks for different initial dark matter profiles, including DM halos with cored profiles and those with extremely cuspy profiles. Cored profiled can result from non-gravitational interactions between DM particles, e.g. in FDM [15, 16] and SIDM [19, 20] models, while extremely cuspy profiles are found in SIDM model when a halo undergoes core collapse [54, 55, 56, 57, 58]. Notably, Penarrubia et al. [51] has found that for DM profiles with different inner slopes, the tidal track is significantly different. We verify such dependence on the inner slope and also investigate the influence of the density slope at larger radii.
@@
We then build improved semi-analytic models (SAMs) for the tidal effects, including tidal stripping and tidal heating, and calibrate these models to the idealized simulations. In previous work [59, 60, 61, 57], similar semi-analytic models have been used to describe the tidal evolution of DM halos initialized with Navarro–Frenk–White (NFW) density profiles [62]. To model the density evolution of a subhalo due to tidal heating, a commonly used approach is to estimate the heating energy rate using the impulse approximation [63]. From the heating energy, the expansion of mass shells in a subhalo can be solved and converted to the change in density profile [59, 61]. In one of our previous studies  [64], we showed that by including the contribution from the second-order perturbation in the heating energy, we can reproduce the tidal track for NFW halos accurately. In this work we will extend these models to other dark matter profiles.
@@ II.1 Initial conditions
At the initial time the subhalo is assumed to be spherically symmetric and has a density profile described by [69, 70]:
@@
To enforce that the subhalo has a finite total mass, we truncate the subhalo density profile at $r>r_{\mathrm{vir}}$ following [71]:
@@
Given the density profile of a subhalo, an N-body realization is generated by sampling particle positions from that density profile. For the initial velocities of particles we assume an isotropic velocity dispersion and use Eddington’s formula [73] to compute the velocity distribution function as a function of radius, taking into account the effects of gravitational softening using the approach of [67]. Velocities are then sampled from that distribution function at the position of each particle. We take a particle mass of $M_{\mathrm{p}}=10^{2}\,\mathrm{M}_{\odot}$ so that the subhalo contains $N\sim 10^{7}$ particles within its virial radius initially. We limit our analysis to the simulation outputs when the subhalo is still resolved by at least $10^{4}$ particles so that its density profile can be well measured. For the choice of softening length, we follow Ref. [49], in which it is suggested that for $N=10^{6}$ a value of $r_{\mathrm{softening}}=0.0003r_{\mathrm{vir,sub}}$ is sufficient to mitigate the numerical artefacts [74]. We rescale the softening length according to the particle number, i.e. $r_{\mathrm{softening}}\propto N_{p}^{1/3}$ [75] and take
@@ II.2 Orbital and tidal evolution
The GADGET-4 code  [76] is used to simulate the evolution of the subhalo within the static host potential. Instead of the TreePM method commonly used in previous simulations, we make use of a new feature in GADGET-4, the so-called fast multipole method (FMM) which has the advantage that the momentum is better conserved. We did not switch on the “PMGRID” option since we find for the simulations we perform the pure FMM method is faster and more accurate. The parameters controlling the accuracy of force computation and time integration are taken to be:
@@
Our tests show that the above values are sufficient for most of the cases we have run (but see Sec. III.1.1 for the exceptional case with an extremely cuspy profile). Details of each parameter can be found in the Gadget-4 documentation  [^2] https://wwwmpa.mpa-garching.mpg.de/gadget4/gadget4_manual.pdf.
@@
where $\eta$ is set by the parameter “ErrTolIntAccuracy”, and $\Delta t_{\rm max}$ is set by “MaxSizeTimestep”. For the extremely cuspy case with $\gamma=1.5$, we find that the results for the tidal track are not converged with the fiducial settings and an extremely small $\eta$ is needed to achieve convergence, so we also test a fixed, global time step in which case all particles are forced to have the same time step size determined by the minimum over $\Delta t_{\mathrm{i}}$ (see Sec. III.1.1).
@@ II.3 Analysis
A reliable self-binding analysis is very important for identifying subhalos and determining their properties (such as the bound mass). Typically, in cosmological simulations, various halo finders such as Rockstar [77] and AHF [78] have been employed in both cosmological and idealized simulations (see, for example [79]). Robust detection and characterization of subhalos in such simulations is challenging—for example, recent studies show that the commonly-used Rockstar halo finder fails to find a non-negligible fraction of subhalos [80, 81]. In the present work we have the advantage that all particles begin as members of the subhalo—we therefore know the starting point which allows us to proceed in a more careful and controlled manner from one snapshot to the next. Our self-binding algorithm is implemented in the Galacticus code [82] to perform this analysis.
@@
    At the beginning of simulation, all particles are assumed to be gravitationally bound to the subhalo. For snapshots at later times, we use the bound/unbound status of particles from the previous snapshot in the first iteration of our algorithm;
@@
    The center-of-mass position ${\mathbf{r}}_{\rm cm}$ and velocity $\mathbf{v}_{\rm cm}$ are determined from all bound particles;
@@
    Repeat steps (ii)–(iv) until the following criterion is satisfied:
@@
This procedure is similar to the one used by van den Bosch & Ogiya [48]. In Ref. [48], the authors use the $5\%$ most-bound particles to determine the center-of-mass properties of the subhalo. In our analysis, we compute ${\mathbf{r}}_{\rm cm}$ and ${\mathbf{v}}_{\rm cm}$ using all bound particles, which in general gives smoother results for $M_{\rm bound}$. Furthermore, our convergence criteria differ slightly from that of Ref. [48], based on the bound mass rather than the center-of-mass position and velocity.
The evolution of a subhalo’s density profile as it orbits within a host potential can be characterized by its $V_{\mathrm{max}}$ and $R_{\mathrm{max}}$. As the subhalo evolves, it is subjected to the tidal force of the host halo. $V_{\rm max}/V_{\rm max,0}$ and $R_{\rm max}/R_{\rm max,0}$ evolve along a so-called called “tidal track” which is largely independent of precisely how mass is stripped away from the subhalos [50, 51, 52, 53]. Here $V_{\rm max,0}$ and $R_{\rm max,0}$ are the initial values of these parameters. While the tidal track is insensitive to the host potential, it does show significant differences for different initial density profiles [51]. Therefore, finding the tidal tracks for different dark matter profiles is crucial for modeling how subhalo evolution in a host potential may differ under various assumptions for the nature of the dark matter particle.
@@ III.1 Tidal tracks
The tidal tracks, $V_{\mathrm{max}}/V_{\mathrm{max},0}$ versus $R_{\mathrm{max}}/R_{\mathrm{max},0}$, for different combinations of $(\alpha,\beta,\gamma)$ are shown in Figs. 1, 2 and 3. First, we keep $\alpha$ and $\beta$ fixed at $1$ and $3$ (the values for an NFW profile), respectively, and change $\gamma$ (the logarithmic slope of density profile at small radii), see Fig. 1. As can be seen, for $\gamma\leq 1$, the results from GADGET-4 are in good agreement with the fitting functions found by [51, dotted] and [53, red dashed, for $\gamma=1$ only]. However, for $\gamma=1.5$ we find that the tidal track obtained from GADGET-4 simulations deviates from the [51] fitting function at late times (i.e. for small values of $V_{\mathrm{max}}/V_{\mathrm{max},0}$ and $R_{\mathrm{max}}/R_{\mathrm{max},0}$). Increasing or decreasing the gravitational softening length by a factor of $3$ slightly changes our results, but can not explain the large deviation from the fitting curve.
Figure 1: Tidal tracks for $(\alpha,\beta)=(1,3)$ and different values of $\gamma$ from N-body simulations using GADGET-4. In the lower right panel, $\eta$ is the parameter ErrTolIntAccuracy in GADGET-4 that controls the time step size. Lines show the fitting functions found of [51, dotted], [53, red dashed, for $\gamma=1$ only], and this work (solid). Note that for the case with $\gamma=0.5$, the last data point, i.e. the one with the lowest $V_{\mathrm{max}}/V_{\mathrm{max},0}$ has been excluded from fitting.
@@
We next increase $\beta$ from $3$ to $4$, which produces a more rapid decrease of density at large radii—results are shown in Fig. 2. Compared to the cases with $\beta=3$ and the same $\gamma$, the cases with $\beta=4$ are less influenced by tidal effects. For example, if we compare the blue circles in Fig. 2 with the blue circles in Fig. 1, both having the same $R_{\mathrm{p}}/R_{\mathrm{a}}$, at the end of the simulation $V_{\mathrm{max}}/V_{\mathrm{max},0}$ and $R_{\mathrm{max}}/R_{\mathrm{max},0}$ change less relative to their initial values than in the case with $\beta=4$. However, the slope of the tidal track in the case with $\beta=4$ is steeper.
Figure 2: Tidal tracks for $(\alpha,\beta)=(1,4)$ and different values of $\gamma$ from N-body simulations using GADGET-4. Solid lines show the fitting functions found in this work.
Finally, we also run two cases with a different value of $\alpha=2$ as shown in Fig. 3. The parameter $\alpha$ controls how smoothly the density profile slope transits from inner value to the value at larger radii. For this value of $\alpha$, we limit our simulations to cored profiles with $\gamma=0$ and $\beta=3,4$. Such density profiles have been found to well describe the observed density profile of Galactic globular clusters that are unaffected by tidal effects [84] [^3] Our parameter $\beta$ corresponds to the parameter $\gamma$ in Ref. [84]..
@@
Table 1: Best-fit parameters for equation (19) for tidal tracks with different initial density profiles.
@@
For the case $\alpha=1$, $\beta=3$ and $\gamma=1$, Ref. [53] proposed a slightly different fitting formula:

### v2

# Tidal evolution of cored and cuspy dark matter halos

@@ Abstract
The internal structure and abundance of dark matter halos and subhalos are powerful probes of the nature of dark matter. In order to compare observations with dark matter models, accurate theoretical predictions of these quantities are needed. We present a fast and accurate method to describe the tidal evolution of subhalos within their parent halo, based on a semi-analytic approach. We first consider idealized $N$-body simulations of subhalos within their host halo, using a generalized mass density profile that describes their properties in a variety of dark matter models at infall, including popular warm, cold, and self-interacting ones. Using these simulations we construct tidal “tracks” for the evolution of subhalos based on their conditions at infall. Second, we use the results of these simulations to build semi-analytic models for tidal effects, including stripping and heating and implement them within the code galacticus. Our semi-analytic models can accurately predict the tidal evolution of both cored and cuspy subhalos, including the bound mass and density profiles, providing a powerful and efficient tool for studying the post-infall properties of subhalos in different dark matter models.
@@ I Introduction
However, while the CDM model has shown excellent agreement with observations on large scales, such as the Cosmic Microwave Background and large-scale structure of galaxy distributions, reproducing some observations on subgalactic scales is challenging within the model [1, 2]. Examples of these challenges include the cusp-core problem [3, 4, 5], the too-big-to fail problem  [6], and diversity problem [7, 8]. To solve these issues, a number of alternative DM models have been proposed, including, e.g. warm dark matter [9, 10, 11, 12], fuzzy dark matter (FDM) [13, 14, 15, 16, 17, 18], self-interacting dark matter (SIDM) [19, 20], and primordial black holes [21, 22, 23].
A particularly powerful probe of the nature of DM is the abundance and internal structure of halos and subhalos. These are the bound hierarchical structures that form as a result of gravity, as the Universe evolves. Alternate dark matter models predict different density profiles and abundance for small DM halos and subhalos. The demographics and internal structure of halos and subhalos already provide some of the most stringent limits on alternative dark matter models, including from observations of Milky Way satellites [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] and strong gravitational lensing [35, 36, 29, 37, 38, 39, 40].
@@
One way to control numerical artifacts is to run idealized simulations, in which case only one subhalo is evolved in the host potential so that one can run high-resolution simulations with manageable computing resource. Using such methods, van den Bosch and Ogiya [49] carried out a detailed study of numerical artifacts in simulations resulting from gravitational softening, discreteness noise, and two-body relaxation. They also derived the requirements for obtaining properly converged results. Using a similar approach, Aguirre-Santaella et al. [50] studied the evolution of subhalos in a realistic Milky Way-like host potential including contributions from the galactic bulge and disk. However, in order to compute statistical properties of subhalos, one needs a model to describe the tidal evolution of subhalos so that one could look at the evolution of a population of subhalos within the context of a cosmological model. In order to achieve this, one can either build analytic models using appropriate approximations, or train non-parametric models such as the one presented in Ref. [51].
@@
In this work, we first run idealized $N$-body simulations to study the evolution of DM halos with different initial density profiles in a host gravitational potential, aiming to include a broad range of dark matter profiles that may be produced by different DM models. Previous studies [52, 53, 54, 55] have shown that, as subhalos evolve in the host, their maximum circular velocity, $V_{\mathrm{max}}$ [^1] The circular velocity is defined as $V=\sqrt{GM(r)/r}$. and the radius at which this maximum is reached, $R_{\mathrm{max}}$, follow a universal “tidal track” for a specific initial density profile. Both $V_{\rm max}/V_{\rm max,0}$ and $R_{\rm max}/R_{\rm max,0}$ are functions of fractional mass remaining in the subhalo and are not sensitive to how mass is stripped from the subhalos. We calculate the tidal tracks for different initial dark matter profiles, including DM halos with cored profiles and those with extremely cuspy profiles. Cored profiled can result from non-gravitational interactions between DM particles, e.g. in FDM [15, 16] and SIDM [19, 20] models, while extremely cuspy profiles are found in SIDM model when a halo undergoes core collapse [56, 57, 58, 59, 60]. Notably, Penarrubia et al. [53] has found that for DM profiles with different inner slopes, the tidal track is significantly different. We verify such dependence on the inner slope and also investigate the influence of the density slope at larger radii.
@@
We then build improved semi-analytic models (SAMs) for the tidal effects, including tidal stripping and tidal heating, and calibrate these models to the idealized simulations. In previous work [61, 62, 63, 59], similar SAMs have been used to describe the tidal evolution of DM halos initialized with Navarro-Frenk-White (NFW) density profiles [64]. To model the density evolution of a subhalo due to tidal heating, a commonly used approach is to estimate the heating energy rate using the impulse approximation [65]. See also Ref. [66] for another approach using the adiabatic limit. From the heating energy, the expansion of mass shells in a subhalo can be solved and converted to the change in density profile [61, 63]. In one of our previous studies  [67], we showed that by including the contribution from the second-order perturbation in the heating energy, we can reproduce the tidal track for NFW halos accurately. In this work we will extend these models to other dark matter profiles.
@@ II.1 Initial conditions
At the initial time the subhalo is assumed to be spherically symmetric and has a density profile described by [72, 73, 74]:
@@
The above parametrization with three free parameters are also known as the “Nuker” model which was first used to describe the surface brightness profile of galactic nuclei [75]. To enforce that the subhalo has a finite total mass, we truncate the subhalo density profile at $r>r_{\mathrm{vir}}$ following [76]
@@
Given the density profile of a subhalo, an $N$-body realization is generated by sampling particle positions from that density profile. For the initial velocities of particles we assume an isotropic velocity dispersion and use Eddington’s formula [78] to compute the velocity distribution function as a function of radius, taking into account the effects of gravitational softening using the approach of [70]. Velocities are then sampled from that distribution function at the position of each particle. We take a particle mass of $M_{\mathrm{p}}=10^{2}\,\mathrm{M}_{\odot}$ so that the subhalo contains $N\sim 10^{7}$ particles within its virial radius initially. We limit our analysis to the simulation outputs when the subhalo is still resolved by at least $10^{4}$ particles so that its density profile can be well measured. For the choice of softening length, we follow Ref. [51], in which it is suggested that for $N=10^{6}$ a value of $r_{\mathrm{softening}}=0.0003r_{\mathrm{vir,sub}}$ is sufficient to mitigate the numerical artifacts [49]. We rescale the softening length according to the particle number, i.e. $r_{\mathrm{softening}}\propto N_{p}^{1/3}$ [79] and take
@@ II.2 Orbital and tidal evolution
The Gadget-4 code  [80] is used to simulate the evolution of the subhalo within the static host potential. Instead of the TreePM method commonly used in previous simulations, we make use of a new feature in Gadget-4, the so-called fast multipole method (FMM) which has the advantage that the momentum is better conserved. We did not switch on the “PMGRID” option since we find for the simulations we perform the pure FMM method is faster and more accurate. The parameters controlling the accuracy of force computation and time integration are taken to be
@@
Our tests show that the above values are sufficient for most of the cases we have run (but see Sec. III.1.1 for the exceptional case with an extremely cuspy profile). Details of each parameter can be found in the Gadget-4 documentation.  [^2] https://wwwmpa.mpa-garching.mpg.de/gadget4/gadget4_manual.pdf
@@
where $\eta$ is set by the parameter “ErrTolIntAccuracy,” and $\Delta t_{\rm max}$ is set by “MaxSizeTimestep.” For the extremely cuspy case with $\gamma=1.5$, we find that the results for the tidal track are not converged with the fiducial settings and an extremely small $\eta$ is needed to achieve convergence, so we also test a fixed, global time step in which case all particles are forced to have the same time step size determined by the minimum over $\Delta t_{\mathrm{i}}$ (see Sec. III.1.1).
@@ II.3 Analysis
A reliable self-binding analysis is very important for identifying subhalos and determining their properties (such as the bound mass). Typically, in cosmological simulations, various halo finders such as Rockstar [81] and AHF [82] have been employed in both cosmological and idealized simulations (see, for example [83]). Robust detection and characterization of subhalos in such simulations is challenging—for example, recent studies show that the commonly used Rockstar halo finder fails to find a non-negligible fraction of subhalos [84, 85]. In the present work we have the advantage that all particles begin as members of the subhalo—we therefore know the starting point which allows us to proceed in a more careful and controlled manner from one snapshot to the next. Our self-binding algorithm is implemented in the Galacticus code [86] to perform this analysis.
@@
    At the beginning of simulation, all particles are assumed to be gravitationally bound to the subhalo. For snapshots at later times, we use the bound/unbound status of particles from the previous snapshot in the first iteration of our algorithm.
@@
    The center-of-mass position ${\mathbf{r}}_{\rm cm}$ and velocity $\mathbf{v}_{\rm cm}$ are determined from all bound particles.
@@
    Repeat steps (2)–(4) until the following criterion is satisfied:
@@
This procedure is similar to the one used by van den Bosch and Ogiya [49]. In Ref. [49], the authors use the $5\%$ most-bound particles to determine the center-of-mass properties of the subhalo. In our analysis, we compute ${\mathbf{r}}_{\rm cm}$ and ${\mathbf{v}}_{\rm cm}$ using all bound particles, which in general gives smoother results for $M_{\rm bound}$. Furthermore, our convergence criteria differ slightly from that of Ref. [49], based on the bound mass rather than the center-of-mass position and velocity.
The evolution of a subhalo’s density profile as it orbits within a host potential can be characterized by its $V_{\mathrm{max}}$ and $R_{\mathrm{max}}$. As the subhalo evolves, it is subjected to the tidal force of the host halo. $V_{\rm max}/V_{\rm max,0}$ and $R_{\rm max}/R_{\rm max,0}$ evolve along a so-called called tidal track which is largely independent of precisely how mass is stripped away from the subhalos [52, 53, 54, 55]. Here $V_{\rm max,0}$ and $R_{\rm max,0}$ are the initial values of these parameters. While the tidal track is insensitive to the host potential, it does show significant differences for different initial density profiles [53]. Therefore, finding the tidal tracks for different dark matter profiles is crucial for modeling how subhalo evolution in a host potential may differ under various assumptions for the nature of the dark matter particle.
@@ III.1 Tidal tracks
The tidal tracks, $V_{\mathrm{max}}/V_{\mathrm{max},0}$ versus $R_{\mathrm{max}}/R_{\mathrm{max},0}$, for different combinations of $(\alpha,\beta,\gamma)$ are shown in Figs. 1, 2 and 3. First, we keep $\alpha$ and $\beta$ fixed at $1$ and $3$ (the values for an NFW profile), respectively, and change $\gamma$ (the logarithmic slope of density profile at small radii), see Fig. 1. As can be seen, for $\gamma\leq 1$, the results from Gadget-4 are in good agreement with the fitting functions found by [53] (dotted) and [55] (red dashed, for $\gamma=1$ only). However, for $\gamma=1.5$ we find that the tidal track obtained from Gadget-4 simulations deviates from the [53] fitting function at late times (i.e. for small values of $V_{\mathrm{max}}/V_{\mathrm{max},0}$ and $R_{\mathrm{max}}/R_{\mathrm{max},0}$). Increasing or decreasing the gravitational softening length by a factor of $3$ slightly changes our results, but can not explain the large deviation from the fitting curve.
Figure 1: Tidal tracks for $(\alpha,\beta)=(1,3)$ and different values of $\gamma$ and pericentric/apocentric ratios, $R_{p}/R_{a}$, from $N$-body simulations using Gadget-4. Lines show the fitting functions found of [53] (dotted), [55] (red dashed, for $\gamma=1$ only), and this work (solid). Note that for the case with $\gamma=0.5$, the last data point, i.e. the one with the lowest $V_{\mathrm{max}}/V_{\mathrm{max},0}$ has been excluded from fitting. For $\gamma=1.5$ (lower right panel), the results from the fiducial runs (circles and triangles) are not converged below $R_{\mathrm{max}}/R_{\mathrm{max},0}\sim 0.22$ (vertical dashed line). Using a global time step size (squares and stars) leads to better converged results, see the discussions in the Sec. III.1.1.
@@
We next increase $\beta$ from $3$ to $4$, which produces a more rapid decrease of density at large radii—results are shown in Fig. 2. Such profiles have been widely used in modeling the density profile of spherical galaxies, see e.g. [72, 88, 89, 90]. Compared to the cases with $\beta=3$ and the same $\gamma$, the cases with $\beta=4$ are less influenced by tidal effects. For example, if we compare the blue circles in Fig. 2 with the blue circles in Fig. 1, both having the same $R_{\mathrm{p}}/R_{\mathrm{a}}$, at the end of the simulation $V_{\mathrm{max}}/V_{\mathrm{max},0}$ and $R_{\mathrm{max}}/R_{\mathrm{max},0}$ change less relative to their initial values than in the case with $\beta=4$. However, the slope of the tidal track in the case with $\beta=4$ is steeper.
Figure 2: Tidal tracks for $(\alpha,\beta)=(1,4)$ and different values of $\gamma$ from $N$-body simulations using Gadget-4. Solid black lines show the fitting functions found in this work. For comparisons, the fitting functions for the same $(\alpha,\gamma)$, but $\beta=3$ are also shown in each panel (light gray lines).
Finally, we also run two cases with a different value of $\alpha=2$ as shown in Fig. 3. The parameter $\alpha$ controls how smoothly the density profile slope transits from inner value to the value at larger radii. For this value of $\alpha$, we limit our simulations to cored profiles with $\gamma=0$ and $\beta=3,4$. Such density profiles have been found to well describe the observed density profile of Galactic globular clusters that are unaffected by tidal effects [91] [^3] Our parameter $\beta$ corresponds to the parameter $\gamma$ in Ref. [91].
@@
Table 1: Best-fit parameters for Eq. (19) for tidal tracks with different initial density profiles.
@@
For the case $\alpha=1$, $\beta=3$, and $\gamma=1$, Ref. [55] proposed a slightly different fitting formula: